# Can we teach language models to gloss endangered languages?

### v1

@@ Abstract
Interlinear glossed text (IGT) is a popular format in language documentation projects, where each morpheme is labeled with a descriptive annotation. Automating the creation of interlinear glossed text can be desirable to reduce annotator effort and maintain consistency across annotated corpora. Prior research (Ginn et al., 2023; Zhao et al., 2020; Moeller and Hulden, 2018) has explored a number of statistical and neural methods for automatically producing IGT.
@@
Michael Ginn  and Mans Hulden  and Alexis Palmer University of Colorado michael.ginn@colorado.edu
@@ 1 Introduction
We study strategies for selecting in-context examples, finding significant impacts to performance. Our best-performing systems outperform transformer model baselines, despite involving no training whatsoever. They still underperform SOTA systems that induce morphological segmentation, but at the same time hold promise for offering a new approach to interlinear glossing for language documentation practitioners.
@@ 2.3 Related Work
Given the cost and difficulty of obtaining IGT data, research has explored methods to scrape it from Latex documents (Schenner and Nordhoff, 2016; Nordhoff and Krämer, 2022) and even images (Round et al., 2020). Finally, work has attempted to standardize IGT conventions and formats, balancing consistency and expressiveness across languages (Lehmann, 1982; Hughes et al., 2003; Nordhoff, 2020; Mortensen et al., 2023).
@@ 3.1 Data
We use the IGT corpora and splits from the 2023 SIGMORPHON Shared Task (Ginn et al., 2023), allowing us to directly compare to several other systems. We use the languages described in Table 1.
@@ 3.4 Prompting
Though the exact prompt varies from experiment to experiment, all runs use the same base prompt, included in Appendix A. In the system prompt, we define the IGT generation task and desired output format and provide additional information such as the language and list of possible glosses. In the user prompt, we provide few-shot examples (if any) and the target example to be glossed.
@@
Figure 2: Morpheme accuracy of LLM-based glossing on Gitksan, Lezgi, Natugu, and Uspanteko, varying the number of provided examples. Reported values are averages over three runs; error bars indicate standard deviation.
@@ 4.1 Experimental Settings
For all experiments, we run two settings, one with just the base task description, and one where we include a list of possible glosses for functional morphemes. We scrape this list of glosses from all of the seen glosses in the training set. We instruct the model to only use these glosses for functional morphemes (while stem morphemes should still be glossed with their translation). We refer to this setting as [+ Glosslist], with an example gloss list in Appendix B.
@@ 4.2 Results
We report results for our languages in Figure 2, with a full table of results provided in Appendix C. Generally, we see that the model has very weak performance in the zero-shot setting, indicating that the model has little knowledge of our chosen languages. In some cases, the zeroshot experiments produce results that are not even in the desired output format.
@@
The average difference is close to 0, well within a standard deviation in all cases, and thus there is little evidence to suggest that including the gloss list meaningfully affects performance. A possible explanation is that since the model has very limited prior knowledge of these languages, providing a list of glosses without any explanation or examples does not provide any useful information.
To investigate whether including a gloss list changes the predictions at all, we measure the adherence percentage. This metric is computed by dividing the number of predicted (functional) glosses that adhere to the gloss list by the total number of predicted glosses. We report the distribution over languages and settings in Figure 5.
@@ 5.1 Experimental Settings
We create silver segmentations using Morfessor and compute the recall metric as described earlier, but using morphemes rather than words. We train the segmentation model use the default parameters on the training data and use Viterbi inference to segment test examples. We use the Morfessor 2.0 library (Virpioja et al., 2013).
@@ 5.2 Results
We experimented with the Max Word Coverage setting, where we consider the recall of the selected set of candidates as whole, rather than individually. We observe minimal benefits, in fact underperforming the Word Recall setting in many cases.
@@ 6 Comparison with SOTA
For the LLM-based method, we select the chrF strategy and test with 30 examples for Gitksan and 100 examples for the other languages. We make some small prompt optimizations described in Appendix A, and raise the temperature to 0.2. We use the following language models:
@@ 6.1 Discussion
We observe that the LLM based glossing strategies outperform a simple transformer in all languages, despite using no training whatsoever and using a small fraction of the training set as examples. Of the LLM models, Gemini performs best on three languages. However, we note that Gemini refuses to produce answers for many examples, which we count as completely wrong. If we omit such examples, Gemini’s performance is even higher, achieving 55.9%, 50.8%, and 63.9% accuracy on Lezgi, Natugu, and Uspanteko respectively.
@@ 7 Conclusion
We evaluate three popular closed-source LLMs, but results may vary across other models. In particular, we have not yet considered open-source, local LLMs due to resource constraints.
@@
Appendix A Prompt Format
Transcription: some text in \$language
Translation: translation of the transcription line in \$metalang
Glosses: the gloss line for the transcribed text
The main prompt is as follows:
Here are some complete glossed examples:
\$fewshot_examples
Please gloss the following example in \$metalang.
Transcription: \$transcription
Translation: \$translation
Appendix B Example Gloss List
@@
Appendix C Full Results
@@
Table 3: Full morpheme accuracy results across languages, selection strategies, and number of examples.

### v3

@@ Abstract
Interlinear glossed text (IGT) is a popular format in language documentation projects, where each morpheme is labeled with a descriptive annotation. Automating the creation of interlinear glossed text would be desirable to reduce annotator effort and maintain consistency across annotated corpora. Prior research (Ginn et al., 2023; Zhao et al., 2020; Moeller and Hulden, 2018) has explored a number of statistical and neural methods for automatically producing IGT.
@@
Michael Ginn1and Mans Hulden2and Alexis Palmer1 1University of Colorado  2New College of Florida michael.ginn@colorado.edu
@@ 1 Introduction
We study strategies for selecting in-context examples, finding significant impacts to performance. Our best-performing systems outperform transformer model baselines, despite involving no training whatsoever. They still underperform SOTA systems that induce morphological segmentation, but at the same time hold promise for offering a new approach to interlinear glossing for language documentation practitioners. Our code is available on Github. [^2] https://github.com/michaelpginn/igt-icl
@@ 2.3 Related Work
Given the cost and difficulty of obtaining IGT data, research has explored methods to scrape it from   documents (Schenner and Nordhoff, 2016; Nordhoff and Krämer, 2022) and even images (Round et al., 2020). Finally, another line of work has attempted to standardize IGT conventions and formats, balancing consistency and expressiveness across languages (Lehmann, 1982; Hughes et al., 2003; Nordhoff, 2020; Mortensen et al., 2023).
@@ 3.1 Data
We use the IGT corpora and splits from the 2023 SIGMORPHON Shared Task (Ginn et al., 2023), allowing us to directly compare several other systems. We use the languages described in Table 1.
@@ 3.4 Prompting
Though the exact prompt varies from experiment to experiment, all runs use the same base prompt.
Transcription: some text in \$language
Translation: translation of the transcription line in \$metalang
Glosses: the gloss line for the transcribed text
The main prompt is as follows:
Here are some complete glossed examples:
\$fewshot_examples
Please gloss the following example in \$metalang.
Transcription: \$transcription
Translation: \$translation
@@
Figure 2: Morpheme accuracy of LLM-based glossing on Gitksan, Lezgi, Natugu, Uspanteko, and [New Image Caption], varying the number of provided examples. Reported values are averages over three runs; error bars indicate standard deviation. In the Base +Glosslist setting, we provide a list of possible glosses in the prompt.
@@ 4.1 Experimental Settings
For all experiments, we run two settings, one with just the base task description, and one where we include a list of possible glosses for functional morphemes. We scrape this list of glosses from all of the seen glosses in the training set. We instruct the model to only use these glosses for functional morphemes (while stem morphemes should still be glossed with their translation). We refer to this setting as [+ Glosslist], with an example gloss list in Appendix A.
@@ 4.2 Results
We report results for our languages in Figure 2, with a full table of results provided in Appendix B. Generally, we see that the model has very weak performance in the zero-shot setting, indicating that the model has little knowledge of our chosen languages. In some cases, the zeroshot experiments produce results that are not even in the desired output format.
@@
The average difference is close to 0, well within a standard deviation in all cases, and thus there is little evidence to suggest that including the gloss list meaningfully affects performance. A possible explanation is that since the model has very limited prior knowledge of these languages, providing a simple list of glosses without any explanation or examples does not provide any useful information.
To investigate whether including a gloss list changes the predictions at all, even if it doesn’t improve glossing performance, we measure the adherence percentage. This metric is computed by dividing the number of predicted (functional) glosses that adhere to the gloss list by the total number of predicted glosses. We report the distribution over languages and settings in Figure 5.
@@ 5.1 Experimental Settings
We create silver segmentations using Morfessor and compute the recall metric as described earlier, but using morphemes rather than words. We train the segmentation model using the default parameters on the training data, and use Viterbi inference to segment test examples. We use the Morfessor 2.0 library (Virpioja et al., 2013).
@@ 5.2 Results
We experimented with the Max Word Coverage setting, where we consider the recall of the selected set of candidates as a whole, rather than individually. We observe minimal benefits, in fact underperforming the Word Recall setting in many cases.
@@ 6 Comparison with SOTA
For the LLM-based method, we select the chrF strategy and test with 30 examples for Gitksan and 100 examples for the other languages. We make some small prompt optimizations described in subsection 3.4, and raise the temperature to 0.2. We use the following language models:
@@
  • Meta’s Llama 3.1 8b parameter model (Dubey et al., 2024), using the 8-bit quantization and the MLX (Hannun et al., 2023) checkpoint.
@@ 6.1 Discussion
We observe that the LLM based glossing strategies outperform a simple transformer in nearly every setting, despite using no training whatsoever and using a small fraction of the training set as examples. Even the Llama 8b parameter model, an open-source model that can be run on a laptop, is competitive.
Of the LLM models, Gemini performs best on three languages. However, we note that Gemini refuses to produce answers for many examples, which we count as completely wrong. If we omit such examples, Gemini’s performance is even higher, achieving 55.9%, 50.8%, and 63.9% accuracy on Lezgi, Natugu, and Uspanteko respectively.
@@ 7 Conclusion
We evaluate three popular closed-source LLMs, and one smaller open-source LLM, but results may vary across other models.
@@
Acknowledgements
We thank the anonymous reviewers for their helpful comments. We also thank the Input Experience team at Apple for their feedback on a preliminary version of this work.
This work was supported by compute credits from a Cohere For AI Research Grant, these grants are designed to support academic partners conducting research with the goal of releasing scientific artifacts and data for good projects.
Parts of this work were supported by the National Science Foundation under Grant No. 2149404, “CAREER: From One Language to Another.” Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation.
Appendix A Example Gloss List
@@
We chose to provide just the list of glosses, without any additional information, to replicate the scenario where there are no additional resources other than glossed examples. Of course, if we had access to a dictionary or grammar reference, providing this information could be beneficial.
Appendix B Full Results
@@
Table 3: Full morpheme accuracy results across languages, selection strategies, and number of examples. +Gloss indicates the gloss list was included in the prompt.